Help! I've been working on this problem for hours I justg dont get it.
2x+11 divided by x^2+4x-5 then 6 divided by x+5 is added to it and the entire equation equals 3 divided by x-1
I've gotten this far (x+5)(x-1) is my LCD. I multiply the numberator and the demominator by this i'm stuck when I get to 2x^2+24x=70 what do I do?
The isbn of my text is 0-13-1886846-2 but the problem is from a quiz

Answer by uma(370)

(2x + 11)/(x^2 + 4x - 5) + 6/(x+5) = 3/(x-1)
==> (2x + 11)/(x+5)(x-1) + 6/(x+5) = 3/(x-1)
Multiplying by the LCD throughout we get,
2x + 11 + 6(x-1) = 3(x+5)
==> 2x + 11 + 6x - 6 = 3x + 15
==> 8x + 5 = 3x + 15
==> 8x + 5 - 3x = 3x + 15 - 3x [adding - 3x to both the sides]
==> 5x + 5 = 15
==> 5x + 5 - 5 = 15 - 5
==> 5x = 10
==> 5x/5 = 10/5
==> x = 2
Good Luck!!!